Use any method to find the integrals.

2. S x2 + 4dxxp zuenare SE
Answer #1

Problem 2:

Formulae to remember

1.    \frac{d}{dx}(tanx)=sec^{2}x

2     .\int sec(x) dx= ln(tan(x)+sec(x))

3.   sec^{2}(x)-tan^{2}(x)=1

2. Wx²+4 dx = I Lets consider x=2 tan (u) u dx = 2Sea (u). du - I l utan (w) +4 )x a Secľudu => I = | asec*(w). 2. Vittan*usSec (u) du - Sec (us. tan(u) + £ 1 sec (u du - Sec(u) tan(us + - ln ( Sectus •[ [sec (x) dx = ln( tan(25+Sec (2))] 2 + tan (ui. Integral changes to where c of in the constant integration,

Problem 3:

Formulae to remember:

1.    \frac{d}{dx}(arctanx)=\frac{1}{1+x^{2}}

2. \int \frac{1}{x}= ln|x|

3. Integration by parts

arc (tants de = I We do integration by parts where Judo= uv-sudu where u= arc(tanx) dv = - = - 1 / 2 I@arctamed :. I = arc taNow solve for Stilitz, die 14x²-x² da all4x²) Es under best - Something - de - - 1 z 2 de 2x In(x) = 1 1 2x de Let u=ltro du=Now substituting everything in I= - arctank & Inlal tc where - 1 1 1 1+x² in the constant of integration